For a given reaction, Delta H = -19.9 kJ/mol and D
Answer #1

G = -RT*lnK

K = exp(-G/(RT))

G = H - TS

if G = 0

then

H - TS = 0

H = TS

H/S = T

T = (-19900)/(-55.5) = 358.558 K

this is the equilbirium temperature

K = exp(-G/(RT))

K =1

by definition